Session: Jun 2011 (session 2)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A. Cattani at Sun May 29 16:08:37 2011) Success:

Torun (K. Borkowski at Mon May 30 06:44:15 2011) Success:
No known problems.

Yebes (Susana Garcia at Mon May 30 08:46:08 2011) Success:
Not known problems

Effelsberg (Uwe Bach at Mon May 30 10:37:10 2011) Minor failures:
LCP lost. Due to an error in the script that produces the procedure files for the DBBC only RCP was recorded. dbbcifa (BBC01-02) and dbbcifb (BBC05-06) both contain RCP data.

Westerbork (Geert Kuper at Mon May 30 17:36:38 2011) Success:
No known problems

Onsala (Michael Lindqvist at Tue May 31 04:31:00 2011) Success:
Some rain, windy. Otherwise no known problems.